# Tree balance in phylogenetic models

### v2

@@ 3 Methods: How to measure the power of balance indices
While there are various ways of comparing tree balance indices, in the present manuscript, we focus on comparing their power to distinguish alternative phylogenetic models from a null model $P^{0}_{n}$. As most previous studies (e.g., [20, 29, 21, 1, 7, 22, 25, 26]) we choose the Yule model as the null model. The Yule model also belongs to several parameterized tree model families, which allows us to observe the power of TSS as we gradually move away from the Yule model by changing the respective parameter(s). The null hypothesis is that the trees have been constructed under the null model.
We use two-tailed testing as a foundation for measuring the power of the tree balance indices. In particular, this implies that the alternative hypothesis is any deviation from the null model, i.e., that trees are more balanced or imbalanced than expected by chance. The level of confidence $\alpha$ is set to $5\%$, i.e., at most $5\%$ of trees constructed under the null model will be incorrectly assessed as non-$P^{0}_{n}$ (first type error) and the sample size $N_{d}$ is set to $10^{3}$. Then, given a set $TSS$ of tree shape statistics, a null model $P^{0}_{n}$, an alternative model $P^{a}_{n}$, and the number of leaves $n$, the testing procedure is as follows:
@@
    Approximate the null distributions of the TSS values: For each $t\in TSS$, sample $N_{d}$ many trees under the null model $P^{0}_{n}$ and compute their TSS values $v^{0}_{t}=(v^{0}_{(t,j)})_{j=1,...,N_{d}}$ (see Figure 2).
@@
    Compute the critical region: Using a level of significance of $\alpha=5\%$, compute the 0.025- as well as the 0.975-quantiles of the $v^{0}_{t}$ values for each $t$ yielding $A^{crit}_{t}=(-\infty,y_{0.025}^{*})\ \cup\ (y_{0.975}^{*},\infty)$ as the critical region (for trees with values in this region the null model is rejected) and $[y_{0.025}^{*},y_{0.975}^{*}]$ as the region of acceptance (the null model is maintained) for each $t$.
@@
    Test trees under alternative model: Construct $N_{d}$ many trees under $P^{a}_{n}$. For each tree $i$ compute the TSS values for each $t$ and set $x_{t,i}=1$ if the null model is rejected and $x_{t,i}=0$ otherwise.
@@
    Compute power: For each $t$ calculate the number $R_{t}$ of these trees for which the null model is rejected. Then, the power of $t$ is the proportion of rejected trees or the average over $x_{t}=(x_{t,i})_{i=1,...,N_{d}}$:
    $\text{power}_{(n,P^{0}_{n},P^{a}_{n})}(t)=\overline{x_{t}}=\frac{1}{N_{d}}\cdot\sum_{i=1}^{N_{d}}{x_{t,i}}=\frac{R_{t}}{N_{d}}.$
@@
    Compute confidence interval: Using the central limit theorem we know that for high $N_{d}$ we approximately have a normal distribution $\overline{x_{t}}\approx N\left(\mu,\frac{\sigma^{2}}{\sqrt{N_{d}}}\right)$ where $\mu$ is the overall expected value and $\sigma^{2}$ the (finite) variance of a random variable $x_{t,l}$. This allows us to estimate the radius of a $(1-\alpha)$%-confidence interval by approximating $\sigma$ with the sample standard deviation $sd_{x_{t}}$ (note that $z_{\alpha}\approx 1.96$ for $\alpha=5$%, i.e., a 95%-confidence interval):
    $\text{radius}_{(n,P^{0}_{n},P^{a}_{n})}(t)=z_{\alpha}\cdot\frac{\sigma}{\sqrt{N_{d}}}\approx 1.96\cdot\frac{\text{sd}_{x_{t}}}{\sqrt{N_{d}}}=1.96\cdot\frac{\sqrt{\frac{1}{N_{d}-1}\cdot\sum_{i=1}^{N_{d}}{(x_{t,i}-\overline{x_{t}})^{2}}}}{\sqrt{N_{d}}}=1.96\cdot\frac{\sqrt{\frac{R_{t}(N_{d}-R_{t})}{N_{d}(N_{d}-1)}}}{\sqrt{N_{d}}}.$
    By using $R_{t}(N_{d}-R_{t})\leq(N_{d}/2)^{2}$ we can calculate an upper limit for the radius, namely $1.96\cdot\sqrt{\frac{1}{4(N_{d}-1)}}$, which is approximately $0.031\approx 3\%$ for $N_{d}=10^{3}$.
Figure 2: Top: Histograms of the exact distributions of the Furnas rank and the Colless index under the Yule model, the PDA model, and the ETM for $n=16$. The Furnas rank is a balance index, i.e., trees with higher values can be considered as more balanced, and the Colless index is an imbalance index, i.e. trees with higher values can be considered as more imbalanced. The uniform distribution of the ETM on $\mathcal{BT}^{\ast}_{n}$ is directly visible as the Furnas rank is also a tree enumeration method, which assigns all trees in $\mathcal{BT}^{\ast}_{n}$ their unique number in $\{1,...,|\mathcal{BT}^{\ast}_{n}|\}$ (here $|\mathcal{BT}^{\ast}_{16}|=10,905$). The dotted lines indicate the 0.025- and the 0.975-quantile under the Yule model.
Bottom: Histograms of the distributions of the Colless index under the Aldous’ $\beta$ splitting model for several choices of $\beta$ (2,000 sampled trees for each $\beta$ in this case) for $n=30$.
@@
Second, we only consider suitable indices, which means that we had to exclude one index and limit the evaluation of another one to small $n$ for practical reasons: The Colijn-Plazotta ranking produces too high values and thus causes computational problems even for small $n$, like $n=14$, and therefore had to be omitted. Furthermore, due to its quadratic runtime, the Furnas rank has only been evaluated for $n=30$.
@@
The above-mentioned procedure is performed for all combinations of $n$, TSS, and alternative models.
In Section 5 we will illustrate how to do similar experiments using any set of tree shape statistics, any set of alternative models, and with regard to any null model with the help of our R software package poweRbal provided alongside this manuscript.

### v3

@@ 2 Preliminaries
While the Yule model is one of the oldest and most basic tree models, it is often used as a null model in phylogenetic analyses. However, there are many more models available to describe different evolutionary settings. In Tables 3 to 7 we briefly review a collection of tree models. We refer the reader to the supplementary material [40] for further background and additional models.
Table 3: List of bipartitioning tree models, which iteratively bipartition sets of $m$ leaves into two subsets of $i$ and $m-i$ leaves, respectively. They are defined by a symmetric probability distribution $q_{m}=(q_{m}(i);i=1,...,m-1)$ of splitting a given set of size $m$ into two subsets of sizes $i$ and $m-i$.
Model
Probability $q_{m}(i)$
for $i\in\{1,2,\ldots,m-1\}$
Description and examples
Yule model [61, 16]
$\frac{1}{m-1}$
Given $m$ odd, then splitting it into sizes $1|(m-1)$, $2|(m-2)$,…, or $\lfloor\frac{m}{2}\rfloor|\lceil\frac{m}{2}\rceil$ is equally probable. If $m$ is even, this also holds true, except for the case where $i=m-1=\frac{m}{2}$, in which case the probability is half of that of the other possibilities. See [21, Equations (7) & (8)].
Aldous’ $\beta$-splitting model, $\beta\geq-2$ [67]
$\frac{1}{a_{m}(\beta)}\frac{\Gamma(\beta+i+1)\Gamma(\beta+m-i+1)}{\Gamma(i+1)\Gamma(m-i+1)}$
with gamma function
[…]
Balance increases with $\beta$. Here, $a_{m}(\beta)$ is a normalizing constant to ensure $\sum_{i=1}^{m-1}q_{m}(i)=1$. Equals Yule for $\beta=0$, PDA for $\beta=-3/2$, and comb for $\beta=-2$.
[…]
Model
Children
$(\lambda_{c_{1}},\lambda_{c_{2}})$
Other
$\lambda_{i}$
Description and examples
Yule model [61, 16] $(\lambda_{0},\lambda_{0})$ $\lambda_{0}$ Pure birth process, constant and equal rates.
[…]
Unequal fert. inheritance (IF-diff), simple biased speciation, $\zeta\geq 1$ [22, p. 1177] $\left(\frac{2\zeta}{\zeta+1}\lambda_{p},\frac{2}{\zeta+1}\lambda_{p}\right)$ $\lambda_{i}$ Factor $\zeta$ has long-time and exponential effect (opposite effect on the two children). Proportion $\lambda_{c_{2}}/\lambda_{c_{1}}=1/\zeta$ with $\lambda_{c_{1}}+\lambda_{c_{2}}=2\lambda_{p}$. Equals Yule for $\zeta=1$.
Biased speciation, $\zeta\in[0,1]$, [26, p. 149] $(\zeta\lambda_{p},(1-\zeta)\lambda_{p})$ $\lambda_{i}$ The parent’s rate is divided between its two children according to the given ratio $\zeta$. Note that $\zeta\in[0,0.5]$ yields the same model.
Trait-inspired models
[…]
$(\lambda_{p}+\varepsilon_{1},$
$\lambda_{p}+\varepsilon_{2})$
$\lambda_{i}$ With $\varepsilon_{1,2}\sim N(0,\sigma^{2})$ normally distributed with standard deviation $\sigma$. Rates $<0$ are set to a small value $>0$.
Trait-based models
Children
$(\lambda_{c_{1}},\lambda_{c_{2}})$
Other
$\lambda_{i}$
Children
$(x_{c_{1}},x_{c_{2}})$
Other
$x_{i}$
Description
[…]
$\left(10^{\log(x_{c_{1}})+\varepsilon_{3}}\right.$,
$\quad\left.10^{\log(x_{c_{2}})+\varepsilon_{4}}\right)$
$\lambda_{i}$
$(x_{p}+\varepsilon_{1},$
$x_{p}+\varepsilon_{2})$
$x_{i}$ Inherited trait influences speciation rate. With $\varepsilon_{1,2}\sim N(0,\sigma^{2}_{x})$ and $\varepsilon_{3,4}\sim N(0,\sigma^{2}_{\lambda})$. Trait values $<0$ are set to a small value $>0$. Note that $\log=\log_{10}$. There is also a bounded version of this model with an upper limit for the trait values, e.g., $x_{i}\in(0,20)$ symmetrical around $x_{0}=10$ ($x_{0}=100$ in [31]). Equals Yule for $\sigma_{\lambda}=\sigma_{x}=0$.
[…]
$\left(10^{\log(x_{c_{1}})+\varepsilon_{3}}\right.$,
$\quad\left.10^{\log(x_{c_{2}})+\varepsilon_{4}}\right)$
$\lambda_{i}$
$\left(10^{\log(x_{p})+\varepsilon_{1},}\right.$
$\left.10^{\log(x_{p})+\varepsilon_{2}}\right)$
$x_{i}$ Inherited trait (with proportional changes) influences speciation rate. With $\varepsilon_{1,2}\sim N(0,\sigma^{2}_{x})$ and $\varepsilon_{3,4}\sim N(0,\sigma^{2}_{\lambda})$. Note that $\log=\log_{10}$. Equals Yule for $\sigma_{\lambda}=\sigma_{x}=0$.
Table 5: List of discrete-time tree growing models with extinction. These models act at the current leaves and modify their speciation and extinction rates. The process starts with a single leaf (the root) with speciation rate $\lambda_{0}$ and extinction rate $\mu_{0}$. At each time step one of the currently existing leaves is chosen for a speciation or extinction event. In case of speciation, it is split into two leaves, $c_{1}$ and $c_{2}$. We denote the resulting rates of $c_{1}$ and $c_{2}$ as $(\lambda_{c_{1}},\lambda_{c_{2}})$ and $(\mu_{c_{1}},\mu_{c_{2}})$, respectively. $\lambda_{e}$ and $\mu_{e}$ denote the rates of the leaf chosen for an extinction event.
Model
Children
$(\lambda_{c_{1}},\lambda_{c_{2}})$
$(\mu_{c_{1}},\mu_{c_{2}})$
Other
$\lambda_{i}$
$\mu_{i}$
Extinct
$\lambda_{e}$
$\mu_{e}$
Description and examples
(Simple/constant-rate) birth-death model, $\lambda_{0}>0$, $\mu_{0}\geq 0$ [61, 2]
$(\lambda_{0},\lambda_{0})$
$(\mu_{0},\mu_{0})$
$\lambda_{0}$
$\mu_{0}$
“Reconstructed process”: Birth-death process, where extinct taxa/leaves are removed (and degree-2 vertices are suppressed). Constant and equal speciation and extinction rates. Equals Yule for $\mu_{0}=0$. Distribution on $\mathcal{BT}^{\ast}_{n}$ is the same as for Yule even for $\mu_{0}>0$ [41, p. 5].
Alternative birth-death model, $\lambda_{0}>0$, $\mu_{0}\geq 0$
$(\lambda_{0},\lambda_{0})$
$(\mu_{0},\mu_{0})$
$\lambda_{i}$
$\mu_{i}$
00
00
Extinct taxa/leaves remain in the tree as “fossils” with speciation and extinction rate 0. Note that, as opposed to other models employing fossils in which fossils are ignored and only extant species are counted, the fossils in this model count towards the total number $n$ of leaves. Equals Yule for $\mu_{0}=0$.
Trait-based models
Density-dependent model [68]
$(\lambda_{0},\lambda_{0})$
$(\mu_{N},\mu_{N})$
$\lambda_{0}$
$\mu_{N}$
$\mu_{N}=\begin{cases}\frac{\lambda_{0}N}{N^{\ast}}&\text{if }N<N^{\ast}\\ \lambda_{0}&\text{if $N\geq N^{\ast}$.}\end{cases}$, where $N$ denotes the number of currently existing lineages, and $N^{\ast}$ is an equilibrium number of lineages. In other words, the speciation rate is constant, but the extinction rate increases linearly with the number of co-existing lineages until the equilibrium $N^{\ast}$ is reached at which point $\mu=\lambda_{0}$. Extinct taxa are removed (reconstructed process).
SSE models state-dependent speciation and extinction models (see, e.g., [69, 70, 71, 72, 73, 74]) Family of models extending birth-death models to account for speciation, extinction, and trait evolution. Each species is associated with a specific trait/state (observed or hidden), and can transition to another state, or speciate (resp. go extinct) with rates depending on its current state.
Table 6: List of tree growing models that do not use rates.
Model Description
Ford’s $\alpha$-model, $\alpha\in[0,1]$ [49, 75] Start with a cherry and add an additional temporary root edge $(\rho_{temp},\rho)$. Then, in each step – let $m$ be the current leaf number – choose any edge, where each of the $m-1$ inner edges has probability $p=\frac{\alpha}{m-\alpha}$ and each of the $m$ pendant edges has probability $p=\frac{1-\alpha}{m-\alpha}$, subdivide it, and add a new pendant edge with a leaf in each step. Do this until reaching $n$ leaves. Last, delete the edge incident to $\rho_{temp}$. Equals Yule for $\alpha=0$, PDA for $\alpha=1/2$, comb for $\alpha=1$.
Uniform model, proportional-to-distinguishable-arrangements (or -types) model (PDA) [17] Start with an unrooted tree with 3 leaves. Then, in each step, choose any edge, subdivide it, and add a new pendant edge with a leaf until reaching $n$ leaves. Last, choose an edge to insert a root [76, p. 8].
Table 7: List of distribution-based tree models. These models generate trees according to some probability distribution.
Model Probabilities Description
Uniform model, proportional-to-distinguishable-arrangements (or -types) model (PDA) [17]
$P_{n}(T)=P_{n}(\mathcal{T})\cdot\left|\mathcal{T}_{T}\right|$
$P_{n}(T)$ $=\frac{n!}{2^{s(T)}\cdot(2n-3)!!}$
All phylogenetic trees $\mathcal{T}\in\mathcal{BT}_{n}$ are equiprobable, i.e., $P_{n}(\mathcal{T})=\frac{1}{|\mathcal{BT}_{n}|}=\frac{1}{(2n-3)!!}$.
Uniform model, equiprobable-types-model (ETM)
$P_{n}(T)=\frac{1}{|\mathcal{BT}^{\ast}_{n}|}=\frac{1}{we(n)}$
All rooted binary trees $T\in\mathcal{BT}^{\ast}_{n}$ are equiprobable.
@@ 3 Methods: How to measure the power of balance indices
While there are various ways of comparing tree balance indices, in the present manuscript, we focus on comparing their power to distinguish alternative phylogenetic models from a null model $P^{0}_{n}$. As most previous studies (e.g., [32, 22, 33, 31, 26, 34, 35, 36]) we choose the Yule model as the null model for the main paper (we present some additional results using other null models in the supplementary material [40]). The Yule model also belongs to several parameterized tree model families, which allows us to observe the power of TSS as we gradually move away from the Yule model by changing the respective parameter(s). The null hypothesis $H_{0}$ is that the trees have been constructed under the null model.
We use two-tailed testing as a foundation for measuring the power of the tree balance indices. In particular, this implies that the alternative hypothesis is any deviation from the null model, i.e., that trees are more balanced or imbalanced than expected by chance. The level of confidence $\alpha$ is set to $5\%$, i.e., at most $5\%$ of trees constructed under the null model will be incorrectly assessed as non-$P^{0}_{n}$ (first type error). Then, given a tree shape statistic $t$, a null model $P^{0}_{n}$, an alternative model $P^{a}_{n}$, and the number of leaves $n$, the testing procedure is as follows:
@@
    Compute or approximate null distribution of the TSS values: If $P^{0}_{n}$ is either the Yule model, the PDA model, or the ETM, we can compute the exact distribution (function) of $t$ under $P^{0}_{n}$ for smaller $n$ (our package poweRbal provides this option for $n\leq 20$) by generating all trees $T\in\mathcal{BT}^{\ast}_{n}$ with $n$ leaves, calculating their TSS values and using the probabilities $P_{Y,n}(T)$ given in Equation (1) for the Yule model, and $P_{\textup{PDA},n}(T)$ and $P_{\textup{ETM},n}(T)$ given in Table 7 for the PDA and ETM models.
    For larger $n$ (or if $P^{0}_{n}$ is not Yule, PDA, or ETM) we approximate the distribution (function) by generating $N_{d}=10^{4}$ to $10^{5}$ trees under the null model $P^{0}_{n}$ and computing their TSS values (see Figure 2).
@@
    Create hypothesis test: The corresponding non-parametric two-tailed test $\Phi$ with small-sample correction, where the null hypothesis that a tree $T\in\mathcal{BT}^{\ast}_{n}$ was generated under the null model is rejected if $\Phi(t(T))=1$ and maintained if $\Phi(t(T))=0$, has the form
    $\Phi(t(T))=\begin{cases}1&\text{if }t(T)\not\in[c_{1},c_{2}]\text{, }\\ 0&\text{if }t(T)\in(c_{1},c_{2})\text{, }\\ 1&\text{with probability $p_{1}$ or $p_{2}$ if $t(T)=c_{1}$ or $c_{2}$, respectively,}\\ 0&\text{with probability $(1-p_{1})$ or $(1-p_{2})$ if $t(T)=c_{1}$ or $c_{2}$, respectively,}\end{cases}$
    for a lower and upper bound $c_{1},c_{2}\in\mathbb{R}$. $c_{1}$ is chosen as the largest TSS value of step (1) with $P(t(T)<c_{1}|H_{0})\leq\frac{\alpha}{2}=2.5\%$ and $c_{2}$ as the smallest TSS value with $P(t(T)>c_{2}|H_{0})\leq\frac{\alpha}{2}=2.5\%$. Furthermore, we do a small-sample correction [77, Equation  21] to ensure that $P(\Phi(t(T))=1|H_{0})$ is as close to $\alpha$ as possible (and not lower, which would make the test too conservative). For this purpose, we compute the probabilities $p_{1}=\frac{\alpha/2-P(t(T)<c_{1}|H_{0})}{P(t(T)=c_{1}|H_{0})}$ and $p_{2}=\frac{\alpha/2-P(t(T)>c_{2}|H_{0})}{P(t(T)=c_{2}|H_{0})}$ to reject the null hypothesis if the TSS value equals the lower or upper bound.
@@
    Test trees under alternative model: Construct $N_{a}=10^{3}$ many trees $(T_{i}^{a})_{i=1,...,N_{a}}$ under $P^{a}_{n}$. For each tree $T_{i}^{a}$ compute $t(T_{i}^{a})$ and $\Phi(t(T_{i}^{a}))$.
@@
    Compute power: Calculate the number $R$ of these trees for which the null model is rejected, i.e, $\Phi(t(T_{i}^{a}))=1$. Then, the power of $t$ is the proportion of rejected trees $\frac{R}{N_{a}}$ or the average over $\Phi^{a}\coloneq(\Phi(t(T_{i}^{a})))_{i=1,...,N_{a}}$:
    $\text{power}_{(n,P^{0}_{n},P^{a}_{n})}(t)=\overline{\Phi^{a}}=\frac{1}{N_{a}}\cdot\sum_{i=1}^{N_{a}}{\Phi(t(T_{i}^{a}))}=\frac{R}{N_{a}}.$
@@
    Compute confidence interval: As the $\Phi(t(T_{i}^{a}))$ are i.i.d. with sample mean $\overline{\Phi^{a}}$, we can use the central limit theorem and conclude that for large $N_{a}$ (we use $N_{a}=10^{3}$) we approximately have a normal distribution $\overline{\Phi^{a}}\approx N\left(\mu,\frac{\sigma^{2}}{N_{a}}\right)$, where $\mu$ is the overall expected value and $\sigma^{2}$ the (finite) variance of a random variable $\Phi(t(T_{i}^{a}))$. This allows us to estimate the radius of a $(1-\alpha)$%-confidence interval by approximating $\sigma$ with the sample standard deviation $sd_{\Phi^{a}}$ (note that $z_{\alpha}\approx 1.96$ for $\alpha=5$%, i.e., a 95%-confidence interval):
    $\text{radius}_{(n,P^{0}_{n},P^{a}_{n})}(t)=z_{\alpha}\cdot\frac{\sigma}{\sqrt{N_{a}}}\approx 1.96\cdot\frac{\text{sd}_{\Phi^{a}}}{\sqrt{N_{a}}}=1.96\cdot\frac{\sqrt{\frac{1}{N_{a}-1}\cdot\sum_{i=1}^{N_{a}}{(\Phi(t(T_{i}^{a}))-\overline{\Phi^{a}})^{2}}}}{\sqrt{N_{a}}}=1.96\cdot\frac{\sqrt{\frac{R(N_{a}-R)}{N_{a}(N_{a}-1)}}}{\sqrt{N_{a}}}.$
    By using $R(N_{a}-R)\leq(N_{a}/2)^{2}$ as $0\leq R\leq N_{a}$ we can calculate an upper limit for the radius, namely $1.96\cdot\sqrt{\frac{1}{4(N_{a}-1)}}$, which is approximately $0.031\approx 3\%$ for $N_{a}=10^{3}$.
Figure 2: Histograms of the exact distributions of (a) the Furnas rank and (b) the Colless index under the Yule model, the PDA model, and the ETM for $n=16$. The Furnas rank is a balance index, i.e., trees with higher values can be considered as more balanced, and the Colless index is an imbalance index, i.e., trees with higher values can be considered as more imbalanced. The uniform distribution of the ETM on $\mathcal{BT}^{\ast}_{n}$ is directly visible as the Furnas rank is also a tree enumeration method, which assigns all trees in $\mathcal{BT}^{\ast}_{n}$ a unique number in $\{1,...,|\mathcal{BT}^{\ast}_{n}|\}$ (here $|\mathcal{BT}^{\ast}_{16}|=10,905$). The dotted lines indicate the 0.025- and the 0.975-quantile under the Yule model. In Panel (c), histograms of the distributions of the Colless index under the Aldous’ $\beta$ splitting model for several choices of $\beta$ (2,000 sampled trees for each $\beta$ in this case) for $n=30$ are shown with overlaid density curves, estimated using Gaussian kernel density estimation.
@@
Second, we only consider suitable indices, which means that we had to exclude one index for practical reasons: The Colijn-Plazotta ranking produces too high values and thus causes computational problems even for small $n$, like $n=14$, and therefore had to be omitted.
@@
Steps (1)–(5) are performed for all combinations of $n\in\{30,100\}$, TSS, and alternative models.
Software
In the supplementary material [40], we illustrate how to do similar experiments using any set of tree shape statistics, any set of alternative models, and with regard to any null model with the help of our R software package poweRbal provided alongside this manuscript. While we give more details in the supplementary material, we emphasize that its usage is straightforward. Deciding on a null model, a list of alternative models, and a leaf number $n$, a power analysis encompassing all existing TSS can be done with simple commands (further details on these commands and its options, e.g., specifying the sample sizes, the testing and correction method, and the significance level, can be found in the supplementary material or in the package manual):
# Get all tss
tss <- getAllTSS(n)
# Perform power analysis
power_data <- powerComp(tss, null_model, alt_models, n)
# Plot results
plot(power_data)